# Informed Initialization for Bayesian Optimization and Active Learning

### v1

@@ Appendix A Experimental Setup
We describe the full experimental setup used in the paper: the design of the Bayesian optimization and active learning loops, the benchmarks used, the compute as well as the licenses of all software and datasets. Our code, including all the benchmarks and plotting to reproduce our results, is available at https://github.com/hipeneurips/HIPE.
@@ A.1 Bayesian Optimization Loop
All experiments were conducted using a standardized pipeline based on BoTorch (Balandat et al., 2020) and GPyTorch (Gardner et al., 2018). We use fully Bayesian Gaussian Process (GP) models, with hyperparameter inference performed via the No-U-Turn Sampler (NUTS )(Hoffman and Gelman, 2014) as implemented in Pyro (Bingham et al., 2018). Unless otherwise specified, we draw 192 burn-in samples followed by 288 hyperparameter samples, retaining every 24th sample for evaluation.
Our GP prior is adapted from Hvarfner et al. (2024) to better suit a fully Bayesian setting. Specifically, we set $\mu_{0}=-0.75$ and $\sigma=0.75$, resulting in $\ell_{d}\sim\mathcal{LN}(0.75+\log(D)/2,0.75)$. The noise standard deviation is modeled as $\sigma_{\varepsilon}\sim\mathcal{LN}(-5.5,0.75)$, and the constant mean parameter follows $c\sim\mathcal{N}(0,0.25)$. These modifications ensure that the means of the priors for $\ell_{d}$ and $\sigma_{\varepsilon}^{2}$ approximately match the modes of the corresponding parameters in the prior proposed by Hvarfner et al. (2024), producing similar hyperparameter values in practice.
@@ A.2 Benchmarks
We use three types of benchmarks: synthetic optimization test functions, surrogate-based hyperparameter optimization tasks from LCBench (Zimmer et al., 2021), and high-dimensional SVM hyperparameter optimization problems. Synthetic functions are standard benchmarks for evaluating active learning and Bayesian optimization under controlled noise and dimensionality. LCBench tasks are GP surrogate models trained on 2,000 evaluations of multi-layer perceptrons (MLPs) on real-world datasets, using a Matern 3/2 kernel; all surrogates are included as part of our code release. The SVM benchmarks follow the setup in Ament et al. (2023a), based on the problem originally introduced in Eriksson and Jankowiak (2021). For these tasks, a Support Vector Regressor (Drucker et al., 1996) is fit to the CTSlice dataset, using a fixed subset of 5,000 data points, with 20% reserved for validation. To vary the dimensionality, an XGBoost model is fit to the original 388-dimensional dataset, and the most important features are retained according to XGBoost’s default feature selection criterion. The objective is to minimize the validation RMSE.
@@
In synthetic functions, additive Gaussian noise is introduced directly to the function evaluations. The LCBench benchmarks rely on pre-trained surrogate models, where the posterior mean is evaluated, and in cases where $\sigma_{\varepsilon}$ is non-zero, additional noise is added to the evaluation of the posterior mean.
A.3 Compute Resources
A.4 Licenses
@@ Appendix C Additional Experiments
C.1 Impact of Batch Size on Predictive Performance
@@
C.2 Bayesian Optimization with Different Batch Sizes
@@
C.3 Active Learning with Different Batch Sizes
@@
C.4 Complementary Hyperparameter Visualization on SVM

### v2

@@ Appendix A Experimental Setup
We describe the full experimental setup used in the paper: the design of the Bayesian optimization and active learning loops, the benchmarks used, the compute as well as the licenses of all software and datasets. The code to run the experiments in the paper, including all the benchmarks and plotting to reproduce our results, is available at https://github.com/hipeneurips/HIPE.
An easy-to-run, tested version of the algorithm and accompanying notebook is available open-source in BoTorch. The HIPE acquisition is found at https://github.com/meta-pytorch/botorch/blob/main/botorch_community/acquisition/bayesian_active_learning.py. The notebook, which demonstrates how to run the algorithm should be run in its indented setting and the expected results, can be found at https://github.com/meta-pytorch/botorch/blob/main/notebooks_community/hyperparameter_informed_predictive_exploration.
@@ A.1 Bayesian Optimization Loop
All experiments were conducted using a standardized pipeline based on BoTorch (balandat2020botorch) and GPyTorch (gardner2018gpytorch). We use fully Bayesian Gaussian Process (GP) models, with hyperparameter inference performed via the No-U-Turn Sampler (NUTS )(JMLR:v15:hoffman14a) as implemented in Pyro (bingham2018pyro). Unless otherwise specified, we draw 192 burn-in samples followed by 288 hyperparameter samples, retaining every 24th sample for evaluation.
Our GP prior is adapted from hvarfner2024vanilla to better suit a fully Bayesian setting. Specifically, we set $\mu_{0}=-0.75$ and $\sigma=0.75$, resulting in $\ell_{d}\sim\mathcal{LN}(0.75+\log(D)/2,0.75)$. The noise standard deviation is modeled as $\sigma_{\varepsilon}\sim\mathcal{LN}(-5.5,0.75)$, and the constant mean parameter follows $c\sim\mathcal{N}(0,0.25)$. These modifications ensure that the means of the priors for $\ell_{d}$ and $\sigma_{\varepsilon}^{2}$ approximately match the modes of the corresponding parameters in the prior proposed by hvarfner2024vanilla, producing similar hyperparameter values in practice.
@@ A.2 Benchmarks
We use three types of benchmarks: synthetic optimization test functions, surrogate-based hyperparameter optimization tasks from LCBench (ZimLin2021a), and high-dimensional SVM hyperparameter optimization problems. Synthetic functions are standard benchmarks for evaluating active learning and Bayesian optimization under controlled noise and dimensionality. LCBench tasks are GP surrogate models trained on 2,000 evaluations of multi-layer perceptrons (MLPs) on real-world datasets, using a Matern 3/2 kernel; all surrogates are included as part of our code release. The SVM benchmarks follow the setup in ament2023unexpected, based on the problem originally introduced in eriksson2021high. For these tasks, a Support Vector Regressor (svr) is fit to the CTSlice dataset, using a fixed subset of 5,000 data points, with 20% reserved for validation. To vary the dimensionality, an XGBoost model is fit to the original 388-dimensional dataset, and the most important features are retained according to XGBoost’s default feature selection criterion. The objective is to minimize the validation RMSE. In synthetic functions, additive Gaussian noise is introduced directly to the function evaluations (see Table 1 for details). The LCBench benchmarks rely on pre-trained surrogate models, where the posterior mean is evaluated, and in cases where $\sigma_{\varepsilon}$ is non-zero, additional noise is added to the evaluation of the posterior mean.
@@
A.3 Licenses
@@ Appendix C Additional Experiments
C.1 Runtime Analysis
We compare acquisition function optimization runtime of HIPE and NIPV, and put that in relation to the runtime of the other components of a BO loop that employs fully Bayesian GP modeling. In Table 2 we demonstrate that while HIPE is substantially slower to optimize than NIPV, its total runtime still amounts to only a minor share of the total runtime of the overall BO run.
Runtime Component (sec) HIPE NIPV Random
Model fit (initial) 21.91 $\pm$ 0.15 21.91 $\pm$ 0.15 N/A
Acquisition Opt. 19.61 $\pm$ 1.12 8.62 $\pm$ 0.37
Model fit (BO loop) 181.79 $\pm$ 7.60 189.90 $\pm$ 5.17 177.33 $\pm$ 7.69
qLogNEI Optimization 4.97 $\pm$ 0.18 4.44 $\pm$ 0.05 4.35 $\pm$ 0.06
C.2 Impact of Batch Size on Predictive Performance
@@
C.3 Bayesian Optimization with Different Batch Sizes
@@
C.4 Active Learning with Different Batch Sizes
@@
C.5 Complementary Hyperparameter Visualization on SVM